In a class of 35 students it is known that 24 of them do arts, 20 do chemistry and 22 do biology. All the students do at least one of the 3 subjects, 3 do all the three subjects while 7 do art and biology. 6 do art and chemistry but not biology and 8 do chemistry and biology. How many of them do chemistry or biology only or arts only?


1
Expert's answer
2021-03-31T17:00:47-0400

In a class of 45 students 


"N(A\cup C\cup B)=45"

It is known that 24 of them do arts, 20 do chemistry and 22 do biology. 

"N(A)=24, N(C)=20, N(B)=22"

3 do all the three subjects


"N(A\cap C\cap B)=3"

7 do art and biology


"N(A\cap B)=7"

6 do art and chemistry but not biology


"N(A\cap C \cap B^C)=6"

8 do chemistry and biology

"N(C\cap B)=8"

Then

"N(A\cap B \cap C^C)=N(A\cap B)-N(A\cap B\cap C)"

"=7-3=4"

"N(B\cap C\cap A^C)=N(B\cap C)-N(A\cap B\cap C)"

"8-3=5"




"N(A\cap B^C\cap C^C)=N(A)-N(A\cap B\cap C^C)"

"-N(A\cap C\cap B^C)-N(A\cap B\cap C)"

"=24-4-6-3=11"


"N(B\cap A^C\cap C^C)=N(B)-N(B\cap A\cap C^C)"

"-N(B\cap C\cap A^C)-N(A\cap B\cap C)"

"=22-4-5-3=10"


"N(C\cap A^C\cap B^C)=N(C)-N(C\cap A\cap B^C)"

"-N(C\cap B\cap A^C)-N(A\cap B\cap C)"

"=20-6-5-3=6"

Check


"11+10+6+4+6+5+3=45"



"N(\text{chemistry or biology only})=N(C\cap B\cap A^C)=5"


"N(\text{arts only })=N(A\cap B^C\cap C^C)=11"